OpenStudy (anonymous):

6x^2 + x = 2 . How ?

OpenStudy (anonymous):

Know the quadratic formula?

OpenStudy (anonymous):

Don't Remember How Its Goes ; Like The Order Of It .

OpenStudy (anonymous):

\[\Large \begin{array}{l} 6{x^2} + x = 2\\ 6{x^2} + x - 2 = 0 \end{array}\] Use the quadratic formula:\[\Large \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]To find the roots.

OpenStudy (anonymous):

^ just rearrange the eq. to make it equal to 0. a = x^2 term b = x term c = number term And then you'll come out with two answers for x

OpenStudy (anonymous):

you can also factorize: (3x+2)(2x-1)=0 then solve...

OpenStudy (anonymous):

Factoring works on this one, but not on all of them ...

OpenStudy (anonymous):

Yeah it's definitely worth knowing how to factorise, but not everything has nice integer factors. \[\Large \begin{array}{l} \frac{{ - 1 \pm \sqrt {{1^2} - (4 \cdot 6 \cdot - 2)} }}{{2 \cdot 6}}\\ = \frac{{ - 1 \pm \sqrt {49} }}{{12}}\\ = \frac{{ - 1 \pm 7}}{{12}}\\ = \frac{1}{2}, - \frac{2}{3} \end{array}\]